# Directional Reasoning Trajectory Change (DRTC): Identifying Critical Trace Segments in Reasoning Models

### v1

@@ 4.6 Adjacency and shuffle controls
For curvature roles, we apply a shuffle control: randomly permuting curvature signatures collapses the real CurvImpact–DRTC association toward zero, while the unshuffled assignments remain consistently positive (Appendix A.5, Fig. 10). This supports the claim that curvature similarity reflects structured intervention-response geometry rather than a clustering artifact.
@@
To test robustness beyond the 24-example cross-architecture slice, we run DRTC on $N=500$ MATH problems (Hendrycks et al., 2021) with R1-Distill-Qwen-1.5B (greedy decoding, fixed seeds, $K=8$; median trace length 2694 tokens, IQR [2131, 3068]). The primary falsification gap persists at scale: learned spans (C8) produce stronger pivot-local directional interventions than matched random spans (C9), with median $\Delta=\mathbb{E}_{k}[|\delta|]_{\textsc{C8}}-\mathbb{E}_{k}[|\delta|]_{\textsc{C9}}=0.409$ (95% CI [0.354, 0.453]) and $\Delta>0$ for 355/500 examples (sign test $p=2.3\times 10^{-21}$). Attribution remains moderately concentrated (median Gini 0.586; median top-1 mass 0.089; median top-5% mass 0.209), and curvature remains complementary rather than directional (median Spearman $\rho(|\mathrm{CurvImpact}|,|\mathrm{DRTC}|)=0.080$; median signed $\rho=-0.039$). Full scaling statistics and curvature diagnostic visualizations are provided in Appendix A.3 and Appendix A.4 (e.g., Fig. 9).
@@ 4.8 Qualitative interpretation
Across all four models, DRTC yields consistent qualitative structure in how directional influence is distributed along reasoning traces. Top-magnitude DRTC chunks typically correspond to strategy-setting constraints or key structural commitments (e.g., introducing an algebraic invariant, applying a functional-iteration identity, or enforcing a conjugacy condition). Masking these segments produces the largest pivot-local redirections, consistent with them anchoring the realized trajectory.
Opposing-sign chunks are not arbitrary: they often align with early orientation text, exploratory detours, or provisional reasoning that is later revised. Masking such segments can increase alignment with the realized direction $g$, indicating interpretable directional tension rather than noise. In parallel, chunks with large $|\mathrm{CurvImpact}|$ frequently coincide with visible phase transitions in the trace (e.g., restatement $\rightarrow$ derivation, manipulation $\rightarrow$ verification). While signed curvature–DRTC correlations remain near zero, curvature magnitude tracks attribution magnitude and stratifies the intensity of local reorientation without serving as a causal score. Detailed per-model case studies appear in Appendix B.
@@ 4.9 Summary: DRTC as a process-level steering map
Taken together, the results support DRTC as a process-level steering map for a single realized rollout. The invariance test (Section 4.1) cleanly separates directional attribution from auxiliary geometry logging, while the random-span falsification (Section 4.4) shows that learned pivots identify interventions that are stronger than matched controls under identical masking logic. The concentration results (Section 4.2) imply that DRTC produces an auditable shortlist of influential context segments rather than a diffuse heatmap, and the pivot-weight ablation (Section 4.3; Appendix A.6) confirms that pivot importance modulates aggregation rather than acting as a cosmetic reweighting. Curvature diagnostics (Section 4.5) provide a complementary view of reorientation intensity without functioning as an attribution score, and adjacency/shuffle controls (Section 4.6) reduce the risk that the observed structure is a trivial proximity artifact. Finally, the scaling study (Section 4.7) shows that the falsification signal and concentration behavior persist beyond the small cross-architecture slice.
@@ 5 Conclusion, limitations, and future work
DRTC measures directional redirection at pivot points rather than outcome-level causal effects on correctness, so steering need not imply answer changes. Attribution currently uses fixed-stride chunks and a fixed pivot budget ($K=8$), which stabilizes comparisons but may miss the appropriate granularity of reasoning; adaptive pivots and variable-resolution or hierarchical chunking are natural extensions. Curvature is used strictly as a logit-space geometric descriptor and does not identify the circuits implementing reorientation, motivating integration with mechanistic tracing. Finally, broader evaluation across domains and decoding strategies, together with joint analysis of trajectory steering and outcome sensitivity, is needed to clarify when local redirections accumulate into correctness changes.
@@ Conclusion.
We introduced Directional Reasoning Trajectory Change (DRTC), a process-causal interpretability framework for long-horizon reasoning under on-policy autoregressive generation. DRTC localizes pivot decision points and applies receiver-side interventions that block information flow from selected earlier chunks only at those pivots, yielding signed per-chunk steering scores defined by directional changes in log-probability space. Across four reasoning models, attribution is consistently sparse and structured, learned pivots induce stronger intervention magnitudes than matched random spans, and curvature provides complementary geometric diagnostics without altering DRTC scores. Overall, DRTC provides a trajectory-level account of how specific context elements steer reasoning, while remaining agnostic to outcome-level causality and circuit identification.
@@
We describe the full methodology in the paper, including pivot discovery signals, intervention definition, scoring, and aggregation, along with model and decoding settings. Our experiments use fixed seeds and export complete per-example artifacts (pivot locations, chunk boundaries/identifiers, intervention effects, and attribution scores) to enable independent auditing. Upon acceptance, we will publicly release the full implementation and evaluation code, including scripts to reproduce all figures and tables, as well as the self-contained interactive reviewer UI used for qualitative inspection. For this submission, the relevant code and artifacts will be provided as supplementary material.
@@
A.2 DRTC concentration across models
@@
A.3 Random-span falsification: magnitude primary, structure secondary
@@
A.4 Curvature diagnostic: geometry stratifies intervention-response types
@@
A.5 Adjacency effects and shuffle control
@@
A.6 Pivot weighting ($u_{k}$) concentration
@@ Appendix C Expanded related work
Table 17 situates DRTC relative to recent frameworks for interpreting multi-step reasoning traces. We group prior work by its primary explanatory target and intervention style, and highlight what is unique about DRTC: (i) explicit pivot discovery to localize decision points, (ii) receiver-side interventions applied only at pivots that hold the realized prefix fixed and avoid generating a new continuation, (iii) signed directional steering in log-probability space, and (iv) curvature used strictly as an intervention-response diagnostic rather than an intrinsic importance score.

### v2

@@ 4.2 Robustness, baselines, curvature calibration, and outcome linkage
Curvature calibration.
Curvature is used only diagnostically: it characterizes the geometry of intervention response rather than serving as an attribution or outcome-causal score. Within traces, curvature-impact and DRTC can co-localize, while cross-example aggregates can be small. A calibration and reconciliation of within-trace vs. cross-example statistics is provided in Appendix A.9, with additional cross-model diagnostic visualizations in Appendix A.10.
Outcome linkage (U1b).
Using graded embedding-interpolation interventions and an outcome-facing metric (teacher-forced $\Delta\log p(\text{gold})$), top-ranked DRTC chunks degrade gold-answer log-probability more than strict quartile-matched random controls on a stability-filtered subset. In the stable subset ($|\mathrm{orig\_margin}|>1.0$), the best regime (cs=4=4, $\alpha=0.2$) achieves median paired difference $0.3398$ with 95% bootstrap CI $[0.0269,1.1851]$ and sign fraction $74\%$ (Appendix A.13). We report outcome linkage under greedy decoding with teacher-forced $\Delta\log p(\text{gold})$; robustness across sampling temperatures/decoding regimes and whether end-to-end answer flips track DRTC ranks are left for future work.
@@
To test robustness beyond the 24-example slice, we run DRTC on $N=500$ MATH problems (Hendrycks et al., 2021) with R1-Distill-Qwen-1.5B (greedy decoding, fixed seeds, $K=8$; median trace length 2694 tokens, IQR [2131, 3068]). Compute scales with the number of pivots and evaluated chunks per example; practical cost and memory profiles are summarized in Appendix A.5. The primary falsification gap persists at scale: learned spans (C8) produce stronger pivot-local interventions than matched random spans (C9), with median $\Delta=\mathbb{E}_{k}[|\delta|]_{\textsc{C8}}-\mathbb{E}_{k}[|\delta|]_{\textsc{C9}}=0.409$ (95% CI [0.354, 0.453]) and $\Delta>0$ for 355/500 examples (sign test $p=2.3\times 10^{-21}$). Attribution remains moderately concentrated (median Gini 0.586; median top-1 mass 0.089; median top-5% mass 0.209). Additional scaling statistics and curvature calibration details are in Appendix A.4 and Appendix A.9, with further diagnostic visualizations in Appendix A.10.
@@ 4.4 Qualitative interpretation
Across models, top-magnitude DRTC chunks typically correspond to strategy-setting constraints or key structural commitments (e.g., introducing an invariant or enforcing a structural condition), while opposing-sign chunks often align with early orientation text or exploratory detours that are later revised. Curvature magnitude frequently coincides with visible phase transitions (e.g., restatement $\rightarrow$ derivation), providing a complementary view of reorientation intensity. Detailed case studies and exported reviewer-auditable artifacts appear in Appendix B.
@@ 4.5 Summary: DRTC as a process-level steering map
Taken together, the results support DRTC as a process-level steering map for a single realized rollout: curvature logging is strictly diagnostic (Section 4.1), influence is concentrated into an auditable shortlist (Appendix A.3), learned pivots outperform matched random spans under identical logic (Figure 3; Appendix A.4), and pivot weighting is non-redundant (Appendix A.12). Robustness diagnostics, baseline comparisons, curvature calibration/diagnostics, and graded outcome linkage are provided in Appendix A.2A.13 (see also Appendix A.10).
@@ 5 Conclusion, limitations, and future work
DRTC primarily measures directional redirection at pivot points; while graded interventions show outcome-facing sensitivity using teacher-forced $\Delta\log p(\text{gold})$ under position-matched controls on a stability-filtered subset, such effects do not guarantee end-to-end correctness changes, and flip-based metrics can saturate under strong edits. Attribution currently uses fixed-stride chunks and a fixed pivot budget ($K=8$), which stabilizes comparisons but may miss the appropriate granularity of reasoning; adaptive pivots and variable-resolution or hierarchical chunking are natural extensions. Curvature is used strictly as a logit-space geometric descriptor and does not identify the circuits implementing reorientation, motivating integration with mechanistic tracing. Finally, broader evaluation across domains and decoding strategies, together with stronger outcome-level benchmarks, is needed to clarify when local redirections accumulate into correctness changes.
@@ Conclusion.
We introduced Directional Reasoning Trajectory Change (DRTC), a process-causal interpretability framework for long-horizon reasoning under on-policy autoregressive generation. DRTC localizes pivot decision points and applies receiver-side interventions that block information flow from selected earlier chunks only at those pivots, yielding signed per-chunk steering scores defined by directional changes in log-probability space. Across four reasoning models, attribution is consistently sparse and structured, learned pivots induce stronger intervention magnitudes than matched random spans, and curvature provides complementary geometric diagnostics without altering DRTC scores. Overall, DRTC provides a trajectory-level account of how specific context elements steer reasoning, and graded outcome linkage indicates that top-ranked chunks can have measurable outcome-facing effects (teacher-forced $\Delta\log p(\text{gold})$) under position-matched controls, though broader domains, decoding regimes, and stronger end-task evaluations remain future work.
@@
We describe the full methodology in the paper, including pivot discovery signals, intervention definition, scoring, and aggregation, along with model and decoding settings. Our experiments use fixed seeds and export complete per-example artifacts (pivot locations, chunk boundaries/identifiers, intervention effects, attribution scores, and outcome-linkage summaries) to enable independent auditing. The code and artifact bundle used to generate the tables and figures in this submission are provided as supplementary material; we will publicly release the full implementation and evaluation code, including scripts to reproduce all figures and tables, as well as the self-contained interactive reviewer UI used for qualitative inspection.
@@
A.2 Robustness to direction estimation and trajectory meander
We evaluate robustness to the global direction estimate $g$ by recomputing chunk rankings under multiple direction-estimation strategies (endpoint difference, PCA, robust PCA, piecewise, piecewise-signed) and meander diagnostics. We summarize per-example Spearman correlations between rankings and report the range of meander ratios observed.
Figure 5: Direction-estimation robustness and meander diagnostics (A2).
Comparison Median $\rho$ Min $\rho$ Notes
PCA vs endpoints 1.00 1.00 24/24 identical
Robust PCA vs endpoints 1.00 1.00 24/24 identical
Piecewise vs endpoints $\approx$1.00 0.50 short-trace edge case
Piecewise-signed vs endpoints $\approx$1.00 0.50 same edge case
Table 1: Ranking invariance across $g$ estimators on the 24-example slice (model: R1-Distill-Qwen-1.5B).
A.3 DRTC concentration across models
@@
A.4 Random-span falsification: magnitude primary, structure secondary
@@
A.5 Computational cost
We report end-to-end computational cost for the DRTC pipeline on the 24-example slice (model: R1-Distill-Qwen-1.5B, $K=8$, greedy decoding, fixed seeds). Costs include pivot discovery, receiver-side interventions, and scoring.
Metric Mean Median Min Max
Wall time per example (s) 11.8 12.28 3.72 17.03
Peak GPU memory (MB) 4835 4864 4164 5333
Table 6: Compute cost summary from main_text_table_cost.csv.
A.6 Masking ablation and implementation details
We ablate receiver-side attention-edge masking by masking subsets of heads (and optionally subsets of layers), and compare resulting rankings against the default all-head, all-layer intervention.
Masking variant $\rho$ vs all-heads (range) Notes
Head subsets (fixed layers) 0.22 – 1.00 agreement can degrade sharply
Table 7: Head-subset masking can substantially alter rankings, motivating the default all-head intervention.
Figure 9: Representative ranking agreement for head-subset masking variants.
A.7 Hyperparameter sensitivity analysis
We evaluate sensitivity of chunk rankings to key heuristic knobs (window size, pivot spacing, gating calibration, and scaling/clipping parameters) by sweeping each knob over 2–3 values and measuring Spearman agreement with the default setting.
Knob Values Median $\rho$ range
window_size 3, 5, 8 0.93 – 1.00
min_pivot_spacing 2, 4, 8 0.90 – 1.00
(others) (see CSV) $\geq 0.96$ typical
Table 8: Sensitivity grid summary (full table in the exported CSV).
Figure 10: Sensitivity heatmap over knob sweeps.
A.8 Baseline attribution comparisons
[…]
Method Median $\rho$ vs DRTC Top-5 overlap Top-10 overlap
[…]
Smooth masking 0.45 0.43 0.82
Chunk occlusion 0.30 0.43 0.67
Activation patching 0.04 0.43 0.67
Table 9: Baseline agreement summary (R1-Distill-Qwen-1.5B, 24-example slice).
Figure 11: Distribution of per-example Spearman agreement between DRTC and baseline methods.
A.9 Curvature calibration and reconciliation
We distinguish (i) within-trace co-localization between chunk-level curvature-impact and DRTC, from (ii) cross-example correlations between per-example summary statistics. These measure different quantities and need not agree.
Statistic (within-trace, N=423) Value
Median $\rho(\mathrm{CurvImpact},\mathrm{DRTC})$ 0.77
Bootstrap 95% CI [0.70, 0.80]
Fraction $p<0.05$ 48%
Table 10: Within-trace curvature-impact co-localization at scale (competition-math scaling run).
Figure 12: Distribution of within-trace Spearman correlations for curvature-impact vs DRTC.
A.10 Curvature diagnostic: geometry stratifies intervention-response types
@@
A.11 Adjacency effects and shuffle control
@@
A.12 Pivot weighting ($u_{k}$) concentration
@@
A.13 Outcome linkage via graded interventions
We evaluate whether top-ranked DRTC chunks have stronger outcome-level effects than position-matched random chunks using graded perturbations and an outcome-facing metric: teacher-forced $\Delta\log p(\text{gold})$. Random controls are strictly quartile-matched to the top chunk’s position bin to remove positional confounds.
Cohort Regime $(cs,\alpha)$ NN Median paired diff 95% CI Sign frac
Stable subset ($|\mathrm{orig\_margin}|>1$) (4, 0.2) 19 0.34 [0.03, 1.19] 0.74
All usable examples (8, 0.7) 22 0.44 [0.04, 1.53] 0.73
Table 15: U1b outcome linkage using teacher-forced $\Delta\log p(\text{gold})$. Paired diff is defined as $\mathbb{E}[\Delta\log p(\text{gold})]_{\text{random}}-\Delta\log p(\text{gold})_{\text{top}}$, so positive values indicate the top chunk is more damaging than matched random.
Figure 17: U1b regime scan over $(cs,\alpha)$ showing discrimination on $\Delta\log p(\text{gold})$.
Figure 18: Best-regime effect sizes by chunk type (top/low/random) with bootstrap CIs.
Figure 19: Distribution of paired differences for $\Delta\log p(\text{gold})$ under strict quartile matching.
@@ Appendix C Expanded related work
Table 24 situates DRTC relative to recent frameworks for interpreting multi-step reasoning traces. We group prior work by its primary explanatory target and intervention style, and highlight what is unique about DRTC: (i) explicit pivot discovery to localize decision points, (ii) receiver-side interventions applied only at pivots that hold the realized prefix fixed and avoid generating a new continuation, (iii) signed directional steering in log-probability space, and (iv) curvature used strictly as an intervention-response diagnostic rather than an intrinsic importance score. We also benchmark DRTC against practical chunk-level attribution baselines (Appendix A.8). Circuit-discovery methods such as ACDC/EAP/CD-T target internal mechanisms at a different granularity, and are best viewed as complementary follow-ups seeded by top-DRTC chunks/pivots.
@@
Chunk-level attribution baselines Li et al. [2017]; Sundararajan et al. [2017]; Fong and Vedaldi [2017]; Zhang and Nanda [2024] Chunk importance for a chosen scalar (e.g., logit/probability/directional score) at a target position Input perturbations (remove/replace chunks), gradient saliency, optimized masking, or activation patching at selected layers/positions Provides non-pivoted chunk attributions that do not explicitly localize decision pivots or enforce pivot-local counterfactuals. DRTC differs by (i) discovering pivots, (ii) applying receiver-side masking only at pivots while holding the realized prefix fixed, and (iii) scoring directional redirection relative to the realized trajectory.